A random variable is uniformly distributed between 100 and 150. a. Draw the density function. b. Find P(X > 110). c. Find P(120 < X < 135). d. Find P(X < 122).


1
Expert's answer
2021-05-07T10:34:18-0400

"f(x) = \dfrac{1}{b-a} = \dfrac{1}{50}"


a.) Probability density function can be drawn as,





b.) P(X > 110)


"= (150-110)\dfrac{1}{50}"


"= \dfrac{4}{5}"


"= 0.8"


c.) P(120 < X < 135)

"= (135-120)\dfrac{1}{50}"


"= \dfrac{15}{20}"


"= 0.3"


d.) P(X < 122)

"= (122-110)\dfrac{1}{50}"

"= 0.24"